OpenStudy (anonymous):

PLEASE HELP ME. THIS IS THE THIRD THREAD AND IVE BEEN AT IT FOR MORE THAN 2 HOURS Find the places where the curve is parallel to the y axis (undefined) x^2+xy+y^2=7

OpenStudy (anonymous):

@aum

OpenStudy (loser66):

take derivative isolate y' solve for denominator of derivative . That is where the curve // y axis

OpenStudy (anonymous):

Ok, I have the derivative

OpenStudy (anonymous):

But I don't know what to do with it

OpenStudy (loser66):

give me

OpenStudy (anonymous):

Its -(2x+y)/(2y+x)

OpenStudy (loser66):

ok, let check \((x^2+xy+y^2)'=7'\\ 2x+y+xy'+2yy'=0\\2x +y+y'(x+2y)=0\) \(y'=\dfrac{-x+2y}{2x+y}\) that y' is slope of the curve, and when the curve //y-axis, y' is undefined y' undefined when denominator =0 that mean 2x -y=0 or y = 2x

OpenStudy (anonymous):

I think that you simplified that incorrectly

OpenStudy (anonymous):

I dont understand what they did here

OpenStudy (loser66):

Can you post the original problem? they do dx/dy, not dy/dx

OpenStudy (anonymous):

x^2+xy+y^2=7

OpenStudy (anonymous):

I do dy/dx and dx/dx too

OpenStudy (loser66):

oh, I got it they let minus sign in numerator, so that they have y = -2x now replace y = -2x into the original one \(x^2 +x(-2x)+4x^2=7\) so that \(3x^2=7\\x = \pm\sqrt{7/3}\) replace back to get y = \(\pm 2\sqrt{7/3}\)

OpenStudy (anonymous):

Ok. What did they do after they got the derivative?

OpenStudy (anonymous):

Going back a little bit. I want to work from there

OpenStudy (loser66):

after getting the derivative, let the denominator =0, solve for either x or y.

OpenStudy (anonymous):

So they set x+2y equal to 0?

OpenStudy (loser66):

yes

OpenStudy (anonymous):

Ok. One sec

OpenStudy (anonymous):

I got y=-2x

OpenStudy (anonymous):

Now what?

OpenStudy (loser66):

replace it into original problem, x^2+2xy +y^2 =7 to find x

OpenStudy (anonymous):

Why?

OpenStudy (loser66):

because at that point (x), the curve //y-axis

OpenStudy (anonymous):

Ok, I got 7

OpenStudy (loser66):

ok, let say at y = -2x, y' is undefined. y' is undefined --> curve // y-axis at the curve, where are the point whose y= -2x? just replace to solve

OpenStudy (anonymous):

Oh ok

OpenStudy (loser66):

how to get 7?

OpenStudy (anonymous):

I didn't get 7, my bad

OpenStudy (anonymous):

I got +-SQRT7/3

OpenStudy (loser66):

yes, that is x, right? now plug back y = -2x to get y

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Oh, great!! Thank you very very much!!!

OpenStudy (loser66):

hihihi:) glad to see you get it